Looking for a solution to this questionCompute the amount to be paid for each of the four separate invoices assuming that all invoices are paid within the discount
Answer #1

1.

Merchandise(Gross) Terms Payment
a $10,000 2/10,n/60 $9,800
b $27,500 1/15,EOM $27,225
c $85,000 1/10,n/30 $84,150
d $22,500 3/15,n/45 $21,825

Explanation;

Amount to be paid = Gross Amount × (1- Discount Percentage)

a) $10,000 ×(1- .02) = $9,800 b) $27,500 × (1-.01) = $27,225 c) $85,000 ×(1-.01) = $84,150 d) $22,500 ×(1-.03) = $21,825